OpenStudy (anonymous):

Which value of x satisfies both 3x + y = 9 and 3x − 5y = 15?

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

3x+y=9 y=9-3x put this value in other equation

OpenStudy (anonymous):

3x-5(9-3x)=15 3x-45+15x=15 18x=15+45=60 x=60/18=10/3

OpenStudy (anonymous):

y=9-3*10/3=9-10=-1

OpenStudy (anonymous):

x=10/3 satisfy

OpenStudy (anonymous):

thanks